# Pan-cancer Histopathology WSI Pre-training with Position-aware Masked Autoencoder

### v2

@@ Abstract
Large-scale pre-training models have promoted the development of histopathology image analysis. However, existing self-supervised methods for histopathology images focus on learning patch features, while there is still a lack of available pre-training models for WSI-level feature learning. In this paper, we propose a novel self-supervised learning framework for pan-cancer WSI-level representation pre-training with the designed position-aware masked autoencoder (PAMA). Meanwhile, we propose the position-aware cross-attention (PACA) module with a kernel reorientation (KRO) strategy and an anchor dropout (AD) mechanism. The KRO strategy can capture the complete semantic structure and eliminate ambiguity in WSIs, and the AD contributes to enhancing the robustness and generalization of the model. We evaluated our method on 6 large-scale datasets from multiple organs for pan-cancer classification tasks. The results have demonstrated the effectiveness of PAMA in generalized and discriminative WSI representation learning and pan-cancer WSI pre-training. The proposed method was also compared with 7 WSI analysis methods. The experimental results have indicated that our proposed PAMA is superior to the state-of-the-art methods. The code and checkpoints are available at https://github.com/WkEEn/PAMA.
@@ 1 Introduction
Over the past few years, Transformer structures have made impressive gains in the field of natural language processing [8]. Subsequently, many recent studies have further facilitated the WSI analysis by taking advantage of the Transformer to capture and aggregate long-range information [9, 10, 11]. High-capacity Transformer models have also promoted the development of self-supervised learning [12, 13]. Self-supervised learning pre-trains a large model on proxy tasks to mine enormous amounts of unlabeled data for potential features and then fine-tunes the model on limited data for specific downstream tasks. The emergence of large-scale models has benefited from the Transformer structure and feature mining of massive data through self-supervised learning, e.g. BERT [12], CLIP [14], SAM [15], and GPT series [16, 17]. There are an increasing number of studies fine-tuning the pre-trained models on histopathology images, which achieved promising performance in various tasks [18, 19].
@@
The efficient utilization of unlabeled data firmly fits the trend of pathology image analysis, since there has been an explosion in the volume of pathology image data with the establishment of large open data projects (e.g. the cancer genome atlas program) and the development of online consultation platforms. In this situation, histopathology image foundation models are established based on self-supervised learning frameworks. Typically, Huang et al [20] applied CLIP [14] for multimodal pathology language-image pre-training (PLIP) based on the public data from medical forums. Similarly, Lu et al [21] pre-trained a large-scale visual-language foundation model using over 1.17 million image-caption pairs for histopathology analysis. Ikezogwo et al [22] created a multimodal histopathology dataset QUILT-1M within 1M paired image-text samples for CLIP pre-training. Nevertheless, the models applied in the above studies were originally designed for natural images and language pre-training. Meanwhile, most of the current self-supervised methods for histopathology images focus on learning features of image patches. The substantial resolution of gigapixel WSIs makes it challenging to build an end-to-end framework for WSI-level representation learning. Currently, there is still a lack of available models that can take full advantage of the abundance of histopathology WSIs.
@@
In this paper, we propose a novel self-supervised learning framework named position-aware masked autoencoder (PAMA) for WSI-level representation learning and pan-cancer pre-training. For the very first time, we propose the slide-level mask image modeling (MIM) proxy task that involves spatial structure to reconstruct WSI representation in feature space. Meanwhile, we embed relative distance and orientation information into slide representation and propose a novel cross-attention module with an orientation dynamic updating strategy and an anchor dropout mechanism. We collected 7 large-scale datasets of multiple organs to evaluate the effectiveness of our proposed framework with slide-level representation learning and multi-organ pre-training and compared it with 6 SOTA WSI analysis methods. The experimental results have demonstrated that PAMA is effective in histopathology WSI pre-training and downstream tasks, including cancer sub-typing and biomarkers prediction.
@@
    We evaluated the proposed method on 6 large-scale datasets containing 8,892 WSIs from multiple organs for multiple diagnostic tasks. The results demonstrate that pan-cancer pre-training facilitates PAMA’s significant progress in fine-grained WSI-level tasks, including biomarkers prediction and cancer sub-typing. Furthermore, PAMA achieves the best performance over the other 7 SOTA methods.
@@ 2.2 Self-supervised learning
Self-supervised learning methods have gained considerable interest in computer vision, frequently concentrating on diverse proxy tasks for pre-training without any manual annotations [31, 32, 33, 34]. The label-free approaches facilitate patch representation learning to release resource consumption from fine-grained annotation. Some works focused on context-based methods, such as predicting pathology image cross-stain, predicting the resolution sequence ordering in WSI, and constructing associations between proximity and feature similarity [35, 36, 37]. Other methods leveraged generative models to build proxy tasks that implicitly learn features by minimizing the reconstruction loss in the pixel space, like SD-MAE [31] and MAE-MIL [32]. More approaches applied contrastive learning to enhance patch feature learning. CTransPath [34] proposed a semantically relevant contrastive learning framework that compares relevance between instances to mine more positive pairs. TransPath [38] used BYOL [39] architecture due to its negative sample independence and proposed a token-aggregating and excitation (TAE) module for capturing more global information. However, these patch-level representation learning methods treated patches as independent entities, thereby destroying the integrity of the semantic information in the WSI. Furthermore, under conditions of limited annotation information, such an approach would yield over-fit the slide-level aggregation model.
HIPT [40] investigated the novel concept of slide-level self-learning, representing a significant challenge. Chen et al [40] constructed a two-stage self-supervised framework in which DINO [41] is utilized to pre-train patches (256×256) and then another DINO is pre-trained for the regions (4096×4096) of WSIs. HIPT leveraged the hierarchical structure inherent in WSIs to construct a multi-level self-supervised learning framework. By doing this, the framework learned high-resolution image representations, enabling it to benefit from the plentiful unlabeled WSIs. This contributes to an increase in the accuracy and robustness of tumor recognition. Nevertheless, the ViT backbone employed for HIPT ignores the structural characteristics of large-scale pathology images. Additionally, the multi-stage pre-training may result in the accumulation of bias and error, reducing the performance of the final model.
@@ 3.2.1 Problem formulation
MAE [46] is a promising paradigm for image representation learning. We introduce masked autoencoder into histopathology slide-level representation learning. Unlike natural scene images, histopathology digital images are scale-varying and semantically complex which is challenging to capture the complete semantic structure and eliminate ambiguity. To combat the limitation, we propose a position-aware structure to construct slide representation. Firstly, patch features of a slide are extracted, which is formulated as $\textbf{X}\in\mathbb{R}^{n_{p}\times d_{f}}$, where $n_{p}$ is the number of patches in the slide and $d_{f}$ is the dimension of the patch feature. Inspired by the way to describe spatial information in KAT [30], multiple anchors are selected by clustering the location coordinates of all patches for profiling local structural semantics. The learnable vectors are assigned for these anchors formulated as $\textbf{K}\in\mathbb{R}^{n_{k}\times d_{f}}$, where $n_{k}=[\frac{n_{p}}{c}]$ is the number of anchors in the slide with $c$ representing the expected number of patches per cluster. Additionally, a polar coordinate system is constructed where each anchor is regarded as a pole. In this system, every patch has explicit relative distance and angle definitions to each anchor. Therefore, we define $\textbf{D}\in\mathbb{N}^{n_{k}\times n_{p}}$ and $\textbf{P}\in\mathbb{N}^{n_{k}\times n_{p}}$ to represent the relative distance matrix and the relative polar angle matrix of a WSI, respectively. Specifically, ${D_{ij}}\in\textbf{D}$ and ${P_{ij}}\in\textbf{P}$ correspondingly denote the relative distance and relative polar angle of the $j$-th patch to the $i$-th anchor. Based on the above, a WSI is formulated as $S=\{\textbf{X},\textbf{K},\textbf{D},\textbf{P}\}$. We leverage the anchor-based and position-aware data structure to represent a WSI, which can adaptively maintain the spatial integrity and semantic enrichment of scale-varying slides across multiple organs.
@@ 3.2.2 PAMA encoder
Our proposed position-aware masked autoencoder is shown in Fig.2(I). Some patch tokens of the original WSI feature are randomly masked with a high ratio (i.e. ratio $r=75\%$) and these tokens including their corresponding spatial information are removed. The remaining ($i.e.$, unmasked) tokens are fed into our encoder. Each encoder block is shown in Fig. 2(I)(c), which is formulated as follows
@@ 3.3 Position-aware cross-attention (PACA)
where $\textbf{W}_{q,k,v}\in\mathbb{R}^{d_{f}\times d_{e}}$ are trainable parameters and $d_{e}$ denotes the dimension of the head output, $\varphi_{d}$ and $\varphi_{p}$ are the transformation functions that respectively map the distance and polar angle to corresponding learnable embedding values, $\sigma$ is the softmax function and $n$ is the index of layer.
@@
             $D^{score}=\mathbf{0}$
@@
                   $\textbf{P}^{(n+1)}_{h,i,j}=\textbf{P}^{(n)}_{h,i,j}-\textbf{P}^{(n)}_{h,i,max}$;// Reorientation.
@@ 3.4 Kernel Reorientation (KRO)
As shown in Fig 3, we illustrate the KRO strategy in detail. Regarding the polar angle matrix $\textbf{P}^{(n)}\in\mathbb{N}^{n_{k}\times n_{p}}$ during the n-th block, the initial polar axis is defined as the horizontal direction for all the anchors. For each anchor, the orientation is equally divided into $N$ bins, e.g., if $N=8$, each bin is a $\frac{\pi}{4}$ sector. During the processing of PACA, an attention score matrix of all anchors to patches formulated as $\textbf{A}^{(n)}\in\mathbb{R}^{n_{k}\times n_{p}}$ is obtained which reflects the contribution from patches to anchors. Based on the matrix $\textbf{A}^{(n)}$, we calculate the attention histogram on the orientation for each anchor by summarizing the attention score of all patches within each orientation bin. Then, the bin with the max score is selected as the new polar main axis, i.e., the reorientated polar axis. Based on the new axis of anchors, we update the polar angle of patches and obtain the updated matrix $\textbf{P}^{(n+1)}$. The detailed algorithm is outlined in Algorithm 1.
@@ 4.1 Datasets
We collected three public large-scale datasets from the cancer genome atlas (TCGA) program and three in-house datasets to evaluate our method, which are introduced as follows:
@@
    Endometrium contains 3,654 in-house WSIs of endometrial pathology including 8 categories, namely well/moderately/low-differentiated endometrioid adenocarcinoma (WDEA/MDEA/LDEA), squamous differentiation endometrioid carcinoma (SDEC), plasmacytoid endometrioid carcinoma (PECA), clear cell endometrioid carcinoma (CCEC), mixed-cell endometrioid adenocarcinoma (MCEA), and tumor-free (Normal).
@@
These datasets consist of 8,892 WSIs from multiple organs and can be used for studies such as cancer sub-typing, molecular status prediction, and gene mutation prediction, which are all slide-level tasks. We randomly divided every dataset into training, validation, and testing subsets according to the ratio of 6:1:3, where the training sets were used for multi-organ pre-training and task-specific fine-tuning, validation sets were used to do early stop, and results on the testing sets were reported for evaluation. We describe the task definitions on the six datasets and the utilization of the data under the multi-organ pre-training strategy as shown in Table 1, where tasks are categorized into in-domain and out-of-domain types based on whether or not the fine-tune data is involved in the pre-training process.
Table 1: Definition and data distribution table for all tasks under the multi-organ pre-train strategy, where we define two task types, In-domain and Out-of-domain, based on whether the fine-tune data participates in pre-train or not.
Tasks Organ Pre-train data Fine-tune data Evaluation data Type
Tumor sub-typing (8 types) Endometrium Training subsets of Endometrium, TCGA-NSCLC, TCGA-RCC and USTC-EGFR Training subset of Endometrium Testing subset of Endometrium In-domain
Tumor sub-typing (Normal/LAUD/LUSC) Lung Training subset of TCGA-NSCLC Testing subset of TCGA-NSCLC
Tumor sub-typing (KIRC/KIRP/KICH) Kidney Training subset of TCGA-RCC Testing subset of TCGA-RCC
EGFR gene mutation sub-typing (4 types) Lung Training subset of USTC-EGFR Testing subset of USTC-EGFR
EGFR gene mutation prediction (Mutation/Wild) Lung Training subset of TCGA-EGFR Testing subset of TCGA-EGFR Out-of-domain
HER2 status grading (0/1+/2+/3+) Breast Training subset of BRCA-HER2 Testing subset of BRCA-HER2
@@
Tasks Strategies DINO [41] PLIP [20]
@@
Tumor sub-typing (8 types) w/o pre-train 38.67 0.837 0.424 38.43 0.801 0.384
@@
Tumor sub-typing (Normal/LAUD/LUSC) w/o pre-train 86.19 0.971 0.884 86.43 0.967 0.862
@@
Tumor sub-typing (KIRC/KIRP/KICH) w/o pre-train 91.72 0.978 0.914 85.25 0.976 0.853
@@
EGFR gene mutation sub-typing (4 types) w/o pre-train 83.03 0.804 0.494 83.63 0.813 0.509
@@
EGFR gene mutation prediction (Mutation/Wild) w/o pre-train 84.54 0.737 0.540 81.53 0.613 0.816
@@
HER2 status grading (0/1+/2+/3+) w/o pre-train 46.34 0.794 0.430 47.51 0.782 0.476
single-organ 50.00(+7.89%) 0.801(+0.88%) 0.482(+12.09%) 48.78(+2.67%) 0.799(+2.17%) 0.479(+0.63%)
multi-organ 52.44(+13.16%) 0.811(+2.14%) 0.505(+17.44%) 50.13(+5.51%) 0.812(+3.83%) 0.488(+2.52%)
@@
Figure 4: Improvement on the long-tailed dataset, where (a) shows the categories distribution of the unbalanced Endometrium dataset, (b), (c), and (d) exhibit the ROC curves of each category without pre-training, with pre-training on the single dataset, and with pre-training on multi-organ datasets using DINO patch features
, respectively.

### v4

@@ Abstract
Large-scale pre-training models have promoted the development of histopathology image analysis. However, existing self-supervised methods for histopathology images primarily focus on learning patch features, while there is a notable gap in the availability of pre-training models specifically designed for WSI-level feature learning. In this paper, we propose a novel self-supervised learning framework for pan-cancer WSI-level representation pre-training with the designed position-aware masked autoencoder (PAMA). Meanwhile, we propose the position-aware cross-attention (PACA) module with a kernel reorientation (KRO) strategy and an anchor dropout (AD) mechanism. The KRO strategy can capture the complete semantic structure and eliminate ambiguity in WSIs, and the AD contributes to enhancing the robustness and generalization of the model. We evaluated our method on 7 large-scale datasets from multiple organs for pan-cancer classification tasks. The results have demonstrated the effectiveness and generalization of PAMA in discriminative WSI representation learning and pan-cancer WSI pre-training. The proposed method was also compared with 8 WSI analysis methods. The experimental results have indicated that our proposed PAMA is superior to the state-of-the-art methods. The code and checkpoints are available at https://github.com/WkEEn/PAMA.
@@ 1 Introduction
Over the past few years, Transformer structures have made impressive gains in the field of natural language processing [8]. Subsequently, many recent studies have further facilitated the WSI analysis by taking advantage of the Transformer to capture and aggregate long-range information [9, 10, 11]. High-capacity Transformer models have also promoted the development of self-supervised learning [12, 13]. Self-supervised learning pre-trains a large model on proxy tasks to mine enormous amounts of unlabeled data for potential features and then fine-tunes the model on limited data for specific downstream tasks. The emergence of large-scale models has benefited from the Transformer structure and feature mining of massive data through self-supervised learning, $e.g.$, BERT [12], CLIP [14], SAM [15], and GPT series [16, 17]. There are an increasing number of studies fine-tuning the pre-trained models on histopathology images, which achieved promising performance in various tasks [18, 19].
@@
The efficient utilization of unlabeled data firmly fits the trend of pathology image analysis, since there has been an explosion in the volume of pathology image data with the establishment of large open data projects ($e.g.$, the cancer genome atlas program) and the development of online consultation platforms. In this situation, histopathology image foundation models are established based on self-supervised learning frameworks. Typically, Huang et al [20] applied CLIP [14] for multimodal pathology language-image pre-training (PLIP) based on the public data from medical forums. Similarly, Lu et al [21] pre-trained a large-scale visual-language foundation model using over 1.17 million image-caption pairs for histopathology analysis. Ikezogwo et al [22] created a multimodal histopathology dataset QUILT-1M within 1M paired image-text samples for CLIP pre-training. Nevertheless, the models applied in the above studies were originally designed for natural images and language pre-training. Meanwhile, most of the current self-supervised methods for histopathology images focus on learning features of image patches. The substantial resolution of gigapixel WSIs makes it challenging to build an end-to-end framework for WSI-level representation learning. Currently, there is still a lack of available models that can take full advantage of the abundance of histopathology WSIs.
@@
In this paper, we propose a novel self-supervised learning framework named position-aware masked autoencoder (PAMA) for WSI-level representation learning and pan-cancer pre-training. For the very first time, we propose the slide-level mask image modeling (MIM) proxy task that involves spatial structure to reconstruct WSI representation in feature space. Meanwhile, we embed relative distance and orientation information into slide representation and propose a novel cross-attention module with an orientation dynamic updating strategy and an anchor dropout mechanism. We collected 7 large-scale datasets of multiple organs to evaluate the effectiveness and generalization of our proposed framework with slide-level representation learning and multi-organ pre-training and compared it with 8 SOTA WSI analysis methods. The experimental results have demonstrated that PAMA is effective in histopathology WSI pre-training and downstream tasks, including cancer sub-typing and biomarkers prediction.
@@
    We evaluated the proposed method on 7 large-scale datasets containing 13,685 WSIs from multiple organs for multiple diagnostic tasks. The results demonstrate that pan-cancer pre-training facilitates PAMA’s significant progress in fine-grained WSI-level tasks, including biomarkers prediction and cancer sub-typing. Furthermore, PAMA achieves the best performance over the other 8 SOTA methods.
@@ 2.2 Self-supervised learning
Self-supervised learning methods have gained considerable interest in computer vision, frequently concentrating on diverse proxy tasks for pre-training without any manual annotations [31, 32, 33, 34]. The label-free approaches facilitate patch representation learning to release resource consumption from fine-grained annotation. Some works focused on context-based methods, such as predicting pathology image cross-stain, predicting the resolution sequence ordering in WSI, and constructing associations between proximity and feature similarity [35, 36, 37]. Other methods leveraged generative models to build proxy tasks that implicitly learn features by minimizing the reconstruction loss in the pixel space, like SD-MAE [31] and MAE-MIL [32]. More approaches applied contrastive learning to enhance patch feature learning. CTransPath [34] proposed a semantically relevant contrastive learning framework that compares relevance between instances to mine more positive pairs. TransPath [38] used BYOL [39] architecture due to its negative sample independence and proposed a token-aggregating and excitation (TAE) module for capturing more global information. Chen et al used DINOv2 [40], a state-of-the-art self-supervised learning method based on student–teacher knowledge distillation for pre-training large ViT architectures, for large-scale visual pre-training on 100,426 histology slides. Vorontsov et al [41] presented a million-image-scale pathology foundation model, Virchow, pre-trained on data from approximately 100,000 patients corresponding to approximately 1.5 million WSIs. There are also many studies introduce multimodal data into self-supervised pre-training for histopathology image representation learning [21, 22]. However, these patch-level representation learning methods treated patches as independent entities, thereby destroying the integrity of the semantic information in the WSI. Furthermore, under conditions of limited annotation information, such an approach would yield over-fit the slide-level aggregation model.
HIPT [42] investigated the novel concept of slide-level self-supervised learning, representing a significant challenge. Chen et al [42] constructed a two-stage self-supervised framework in which DINO [43] is utilized to pre-train patches (256×256) and then another DINO is pre-trained for the regions (4096×4096) of WSIs. HIPT leveraged the hierarchical structure inherent in WSIs to construct a multi-level self-supervised learning framework. By doing this, the framework learned high-resolution image representations, enabling it to benefit from the plentiful unlabeled WSIs. This contributes to an increase in the accuracy and robustness of tumor recognition. Recently, Xu et al [44] proposed Prov-GigaPath, a slide-level representation leaning framework pre-trained on 171,189 slides originated from more than 30,000 patients covering 31 major tissue types. Nevertheless, the ViT backbone employed for HIPT ignores the structural characteristics of large-scale pathology images. Additionally, the multi-stage pre-training may result in the accumulation of bias and error, reducing the performance of the final model. Prov-GigaPath applies LongNet [45] as a slide aggregator, leveraging its design for extremely long sequences. However, Prov-GigaPath does not account for the unique characteristics of WSIs, particularly their spatial structure.
@@ 3.2.1 Problem formulation
MAE [50] is a promising paradigm for image representation learning. We introduce masked autoencoder into histopathology slide-level representation learning. Unlike natural scene images, histopathology digital images are scale-varying and semantically complex which is challenging to capture the complete semantic structure and eliminate ambiguity. To combat the limitation, we propose a position-aware structure to construct slide representation. Firstly, patch features of a slide are extracted, which is formulated as $\textbf{X}\in\mathbb{R}^{n_{p}\times d_{f}}$, where $n_{p}$ is the number of patches in the slide and $d_{f}$ is the dimension of the patch feature. Inspired by the way to describe spatial information in KAT [30], multiple anchors are selected by clustering the location coordinates of all patches for profiling local structural semantics. The learnable vectors are assigned for these anchors formulated as $\textbf{K}\in\mathbb{R}^{n_{k}\times d_{f}}$, where $n_{k}=[\frac{n_{p}}{c}]$ is the number of anchors in the slide with $c$ representing the expected number of patches per cluster. Additionally, a polar coordinate system is constructed where each anchor is regarded as a pole. In this system, every patch has explicit relative distance and angle definitions to each anchor. Therefore, we define $\textbf{D}\in\mathbb{N}^{n_{k}\times n_{p}}$ and $\textbf{P}\in\mathbb{N}^{n_{k}\times n_{p}}$ to represent the relative distance matrix and the relative polar angle matrix of a WSI, respectively. We equate the polar and distance values into bins to ensure adaptation to scale-varying slides, so the inputs are discrete integers, similar to the positional index of each token in Transformer[8]. Specifically, ${D_{ij}}\in\textbf{D}$ and ${P_{ij}}\in\textbf{P}$ correspondingly denote the relative distance and relative polar angle of the $j$-th patch to the $i$-th anchor. Based on the above, a WSI is formulated as $S=\{\textbf{X},\textbf{K},\textbf{D},\textbf{P}\}$. We leverage the anchor-based and position-aware data structure to represent a WSI, which can adaptively maintain the spatial integrity and semantic enrichment of scale-varying slides across multiple organs.
@@ 3.2.2 PAMA encoder
Our proposed position-aware masked autoencoder is shown in Fig.2(I). Some patch tokens of the original WSI feature are randomly masked with a high ratio ($i.e.$, ratio $r=75\%$) and these tokens including their corresponding spatial information are removed. The remaining ($i.e.$, unmasked) tokens are fed into our encoder. Each encoder block is shown in Fig. 2(I)(c), which is formulated as follows
@@ 3.3 Position-aware cross-attention (PACA)
where $\textbf{W}_{q,k,v}\in\mathbb{R}^{d_{f}\times d_{e}}$ are trainable parameters and $d_{e}$ denotes the dimension of the head output, $\varphi_{d}$ and $\varphi_{p}$ are the transformation functions that respectively map the distance and polar angle to corresponding learnable embedding values, $\sigma$ is the softmax function and $n$ is the index of layer. We apply two transformation functions to embed polar and distance into vectors, respectively, to ensure that the positional information is continuous and trainable. We add position embeddings as bias in softmax function to effectively facilitate the module to capture global information, drawing inspiration from the Graphormer [51].
@@
$N$: The number of orientation bins assigned to each anchor;
@@
             Initialize $D^{score}$ with $\mathbf{0}$
@@
                   $\textbf{P}^{(n+1)}_{h,i,j}=(\textbf{P}^{(n)}_{h,i,j}-\textbf{P}^{(n)}_{h,i,max})\ mod\ N$; // Reorientation.
@@ 3.4 Kernel Reorientation (KRO)
As shown in Fig 3, we illustrate the KRO strategy in detail. Regarding the polar angle matrix $\textbf{P}^{(n)}\in\mathbb{N}^{n_{k}\times n_{p}}$ during the n-th block, the initial polar axis is defined as the horizontal direction for all the anchors. For each anchor, the orientation is divided into $N$ equal bins. For example, if $N=8$, each bin corresponds to a $\frac{\pi}{4}$ sector. During the processing of PACA, an attention score matrix of all anchors to patches formulated as $\textbf{A}^{(n)}\in\mathbb{R}^{n_{k}\times n_{p}}$ is obtained which reflects the contribution from patches to anchors. Based on the matrix $\textbf{A}^{(n)}$, we calculate the attention histogram on the orientation for each anchor by summarizing the attention score of all patches within each orientation bin. Then, the bin with the max score is selected as the new polar main axis, $i.e.$, the reorientated polar axis. Based on the new axis of anchors, we update the polar angle of patches and obtain the updated matrix $\textbf{P}^{(n+1)}$. The detailed algorithm is outlined in Algorithm 1.
@@ 4.1 Datasets
We collected four public large-scale datasets from the cancer genome atlas (TCGA) program and three in-house datasets to evaluate our method, which are introduced as follows:
@@
    Endometrium-3k contains 3,654 in-house WSIs of endometrial pathology including 8 categories, namely well/moderately/low-differentiated endometrioid adenocarcinoma (WDEA/MDEA/LDEA), squamous differentiation endometrioid carcinoma (SDEC), plasmacytoid endometrioid carcinoma (PECA), clear cell endometrioid carcinoma (CCEC), mixed-cell endometrioid adenocarcinoma (MCEA), and tumor-free (Normal).
@@
  • $\bullet$
    TCGA pan-cancer dataset contains 4,793 unlabeled WSIs containing 10 cancer types from 7 primary sites as shown in Table 1, which is collected from TCGA program designedly for evaluation of generalization for out-of-domain pre-training.
These datasets, except for TCGA pan-cancer dataset, consist of 8,892 WSIs from multiple organs and can be used for studies such as cancer sub-typing, molecular status prediction, and gene mutation prediction, which are all slide-level tasks. We randomly divided every dataset into training, validation, and testing subsets according to the ratio of 6:1:3, where the training sets were used for multi-organ pre-training and task-specific fine-tuning, validation sets were used to do early stop, and results on the testing sets were reported for evaluation. We describe the task definitions on these datasets and the utilization of the data under the multi-organ pre-training strategy as shown in Table 2, where tasks are categorized into in-domain and out-of-domain conditions based on whether or not the fine-tuning data is involved in the pre-training process.
Table 1: Detailed data distribution of TCGA pan-cancer dataset.
Dataset Primary Site Number of WSIs
TCGA-BRCA Breast 1121
TCGA-CESC Gynecology 278
TCGA-BLCA Urinary 457
TCGA-PAAD Liver 205
TCGA-COAD Gastrointestinal 441
TCGA-READ 158
TCGA-STAD 400
TCGA-PRAD Prostate 449
TCGA-GBM Brain 816
TCGA-HNSC Head and Neck 468
Table 2: Definition and data distribution table for all tasks under the multi-organ pre-train strategy, where we define two conditions, In-domain and Out-of-domain, based on whether the fine-tune dataset participates in pre-train or not.
Pre-training condition In-domain Out-of-domain
Pre-training data Training subsets of Endometrium-3k, TCGA-NSCLC, TCGA-RCC,USTC-EGFR, and TCGA-EGFR TCGA pan-cancer dataset
Downstream Tasks
Organ Endometrium Lung Kidney Lung Lung Endometrium Lung Kidney Lung Lung Breast
Fine-tuning data Training subset of Training subset of
Endometrium-3k TCGA-NSCLC TCGA-RCC USTC-EGFR TCGA-EGFR Endometrium-3k TCGA-NSCLC TCGA-RCC USTC-EGFR TCGA-EGFR BRCA-HER2
Evaluation data Testing subset of Testing subset of
Endometrium-3k TCGA-NSCLC TCGA-RCC USTC-EGFR TCGA-EGFR Endometrium-3k TCGA-NSCLC TCGA-RCC USTC-EGFR TCGA-EGFR BRCA-HER2
@@
Datasets Strategies DINO [43] PLIP [20]
@@
Endometrium-3k w/o pre-train 38.67 0.837 0.424 38.43 0.801 0.384
@@
TCGA-NSCLC w/o pre-train 86.19 0.971 0.884 86.43 0.967 0.862
@@
TCGA-RCC w/o pre-train 91.72 0.978 0.914 85.25 0.976 0.853
@@
USTC-EGFR w/o pre-train 83.03 0.804 0.494 83.63 0.813 0.509
@@
TCGA-EGFR w/o pre-train 84.54 0.737 0.540 81.53 0.613 0.816
@@
Figure 4: Improvement on the long-tailed dataset, where (a) shows the categories distribution of the unbalanced Endometrium-3k dataset, (b), (c), and (d) exhibit the ROC curves of each category without pre-training, with pre-training on the single dataset, and with pre-training on multi-organ datasets using DINO patch features, respectively.